A real function f defined on an interval [a,b] with a<c<b where c is a point of the interval, is said to be differentiable at the point x=c if

1
Expert's answer
2022-03-14T19:05:44-0400

A function f is said to be differentiable at x = c, c "\in" (a,b) if both left hand derivative and right hand derivative at x = c exist finitely and they are equal also. That means

(i) "\lim_{x\to c^{-}}\frac{f(x)-f(c)}{x-c}" exists finitely

(ii) "\lim_{x\to c^{+}}\frac{f(x)-f(c)}{x-c}" exists finitely

and (iii)

"\lim_{x\to c^{-}}\frac{f(x)-f(c)}{x-c}"="\lim_{x\to c^{+}}\frac{f(x)-f(c)}{x-c}"